# On the metric property of quantum Wasserstein divergences

### v1

@@ Abstract.
Quantum Wasserstein divergences are modified versions of quantum Wasserstein distances defined by channels, and they are conjectured to be genuine metrics on quantum state spaces by De Palma and Trevisan. We prove triangle inequality for quantum Wasserstein divergences for any finite-dimensional quantum system and any quadratic cost operator under the assumption that a particular state involved is pure. We also provide strong numerical evidence suggesting that the triangle inequality holds in general, for an arbitrary choice of states.
@@
- 5 Appendix: the code used for the numerical study
@@ 1.1. Motivation and main result
For every finite-level quantum system and for every finite collection of observable quantities, the corresponding quadratic quantum Wasserstein divergence satisfies
for any triplet of states assuming that is pure or both and are pure. Moreover, numerical results strongly indicate that (1) holds for any triplet of states without any further assumptions.
@@ 1.2. Basic notions, notation
In quantum mechanics, the state of a particle moving in is described by a wave function of unit norm, or more generally, by a normalized, positive, trace-class operator on Measurable physical quantities correspond to (possibly unbounded) self-adjoint operators on The spectrum of such an operator is precisely the collection of all possible outcomes of a quantum measurement. In the sequel we denote by the real linear space of self-adjoint operators on , and stands for the set of states, that is, the set of positive operators on with unit trace. When measuring an observable quantity on a quantum system being in the state the probability of the outcome lying in an interval is where is the spectral measure of Consequently, a quantum state encapsulates a bunch of classical probability distributions, each corresponding to a physical quantity we are interested in. As the conceptual challenges of quantum optimal transport are present in the finite-dimensional setting, we assume from now on that
@@
That is, a coupling of and is a state on such that
@@
for all Note the clear analogy of the above definition of quantum couplings with the definition of classical couplings that can be rephrased as follows: is a coupling of and if
and
@@
Let us jump back to (6) and note that if the state of our composite quantum system is described by then
Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of and with respect to the measurable quantities is given by
There is a correspondence between quantum couplings of the states (in the sense of (7)), and quantum channels (that is, completely positive, trace-preserving linear maps) defined on the support of that map to . This one-by-one relation is constructive and explicit, see Subsection 2.2. and especially Proposition 2 of [17] for further details. A consequence of this explicit relation — which we will heavily use in the sequel — is that the cost of a coupling can be expressed in terms of the corresponding channel as follows:
See Subsection 5.2. of [18] for the justification of the above relation. Consequently, the quadratic quantum Wasserstein distances are obtained by an optimization over quantum channels:
An important feature of the quadratic Wasserstein distances is that the distance of a state from itself (which may be positive) is always realized by the identity channel — see [17, Corollary 1]. In addition to that, the following concavity-like result is true for any choice of and any :
@@
is a true metric for all cost operator of the form
— up to some non-degeneracy assumptions on the ’s generating to ensure the definiteness of that is, that only if
The fact that the self-distance is realized by the identity channel for every [17, Corollary 1] and (12) together imply that
We denote by the set of rank-one ortho-projections on that is, the set of pure states.
Remark 1.
If either or is pure then the only coupling is the tensor product and the only corresponding channel is , with
@@ 2. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 1
If is pure or both and are pure, then by by Remark 1 and (16) we have
@@
If and are self-adjoint, and , then by the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product on and , and hence we get the following upper bound for the right-hand side (RHS) of (21):
@@
where we used again the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product for This completes the proof of (21) and hence that of the Theorem. ∎
@@ 3. Numerical evidence for the triangle inequality for generic triplets of states
In the first example, we considered and chose the deterministic states and We singled out the section of the Bloch ball, that is, we took ’s of the form
@@ 4.1. Triangle inequality for qubits — the case of the symmetric cost
In this section we stick to and we choose the cost operator (15) to be

### v2

@@ Abstract.
Quantum Wasserstein divergences are modified versions of quantum Wasserstein distances defined by channels, and they are conjectured to be genuine metrics on quantum state spaces by De Palma and Trevisan. We prove triangle inequality for quantum Wasserstein divergences for every quantum system described by a separable Hilbert space and any quadratic cost operator under the assumption that a particular state involved is pure, and all the states have finite energy. We also provide strong numerical evidence suggesting that the triangle inequality holds in general, for an arbitrary choice of states.
@@
- 5 Applications
- 6 Appendix: the code used for the numerical study
@@ 1.1. Motivation and main result
For every quantum system described by a separable Hilbert space and for every finite collection of observable quantities, the corresponding quadratic quantum Wasserstein divergence satisfies
for any triplet of states assuming that is pure or both and are pure, and all the states involved are of finite energy. Moreover, numerical results strongly indicate that (1) holds for any triplet of states without any further assumptions.
@@ 1.2. Basic notions, notation
In quantum mechanics, the state of a particle moving in is described by a wave function of unit norm, or more generally, by a normalized, positive, trace-class operator on Measurable physical quantities correspond to (possibly unbounded) self-adjoint operators on The spectrum of such an operator is precisely the collection of all possible outcomes of a quantum measurement. In the sequel, we denote by the set of self-adjoint but not necessarily bounded operators on , and stands for the set of states, that is, the set of positive trace-class operators on with unit trace. The space of all bounded operators on is denoted by and we recall that the collection of trace-class operators on is denoted by and defied by Similarly, stands for the set of Hilbert-Schmidt operators defined by When measuring an observable quantity on a quantum system being in the state the probability of the outcome lying in an interval is where is the spectral measure of Consequently, a quantum state encapsulates a bunch of classical probability distributions, each corresponding to a physical quantity we are interested in.
@@
where the transpose of a linear operator acting on is a linear operator on the dual space defined by the identity where and That is, a coupling of and is a state on such that
@@
for all bounded Note the clear analogy of the above definition of quantum couplings with the definition of classical couplings that can be rephrased as follows: is a coupling of and if
@@
Let us jump back to (6) and note that by Born’s rule on quantum measurement, if the state of our composite quantum system is described by then
Here we used the convention that the last line of (9) is defined to be if there is an eigenvector of outside the domain of for some Therefore, in view of (6), (7), and (9), the quadratic Wasserstein distance of and with respect to the measurable quantities is given by
We recall (see [18, Definition 6]) that the energy of a state with respect to the observable is given by where is the spectral decomposition of and if for some The energy of with respect to the collection of observables is defined by Recall that classical quadratic Wasserstein spaces consist of probability measures with finite second moment. As the natural quantum analogs of them are states with finite energy, we restrict our attention to such quantum states in the sequel.
By [18, Proposition 3], if the states have finite energy, then any quantum coupling has finite cost. Moreover, both and are trace-class for every – see [18, Lemma 3]. Consequently, by the definition of Hilbert-Schmidt and trace-class operators, and are Hilbert-Schmidt, and so are and as taking the adjoint is an involution of Note furthermore that both and are Hilbert-Schmidt by definition, and hence the operators are trace-class as they are products of two Hilbert-Schmidt operators.
A prominent coupling of a state with itself is the canonical purification which is the rank-one projection corresponding to the unit vector obtained from by the canonical isomorphism between and
An important feature of the quadratic Wasserstein distances is that the distance of a state from itself (which may be positive) is always realized by the canonical purification — see [18, Corollary 1]. That is,
We have seen that the finite energy condition on implies that both and are Hilbert-Schmidt operators, and hence not only but also and and are trace-class operators. We note that taking the adjoint leaves both the Hilbert-Schmidt norm and the trace invariant, and hence and Consequently,
Moreover, the following concavity-like result is true for any choice of and for any with finite energy:
@@
is a true metric for all finite collections of observables on the set of those states on that have finite energy with respect to — up to some non-degeneracy assumptions on the ’s to ensure the definiteness of that is, that only if
Let us denote by the set of rank-one ortho-projections on that is, the set of pure states. Note that if either or is a pure state, then the only coupling of them is the tensor product, that is, Therefore, by (10), the quadratic quantum Wasserstein distance has the following explicit form in this special case:
Here we used that if and that for and that the operators are trace-class.
@@ 2. Triangle inequality for quantum Wasserstein divergences — the proof of Theorem 1
Let be a separable Hilbert space, and let be an arbitrary finite collection of observable quantities, and let be the corresponding quadratic quantum Wasserstein divergence defined by (13) and (10). Let and assume that or both and are in Moreover, assume that and have finite energy with respect to Then the triangle inequality
holds true.
@@
If is pure or both and are pure, then by (14) and (11) we have
@@
If is a state having finite energy with respect to the observable then is a Hilbert-Schmidt operator. Indeed, let be the spectral resolution of and let us compute the trace of the positive operator by
Now let us note that this is precisely the trace of that we already have shown to be a trace-class operator. This latter statement can be checked by the direct computation
So both and are Hilbert-Schmidt operators, and by the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product of them we get
where we used that both and are trace-class and hence their traces coincide, and hence we get the following upper bound for the right-hand side (RHS) of (19):
@@
where we used again the Cauchy-Schwarz inequality for the Hilbert-Schmidt inner product
for and This completes the proof of (19) and hence that of the Theorem.
@@ 3. Numerical evidence for the triangle inequality for generic triplets of states
In the first example, we considered and chose the deterministic states
We singled out the section of the Bloch ball, that is, we took ’s of the form
@@ 4.1. Triangle inequality for qubits — the case of the symmetric cost
In this section we stick to and we choose the cost operator to be
@@ 4.2. Lower estimation of the cost function
5. Applications
The aim of this section is to highlight a direct application of our main result in quantum complexity theory (Subsection 5.1), and to present a brief description of the role of quantum optimal transport distances in mathematical physics (Subsections 5.2 and 5.3).
5.1. Wasserstein complexity
We may define the Wasserstein complexity of quantum channels relying on quantum Wasserstein divergences. Given a finite collections of observables , let us define the Wasserstein complexity of a channel by
In [39] the quantum Wasserstein complexity is introduced in a similar way, using the quantum generalization of the Hamming-Wasserstein classical metric. A connection between this complexity and the circuit cost of unitary channels is provided, the latter being bounded from below by the former. The argument of this bound relies heavily on the triangle inequality of the Wasserstein distance used. We suggest now that (37) may have a similar application in light of the triangle inequality proposed in this paper.
To justify our suggestion we now prove a few useful porperties of (37) that is desirable of a complexity quantity. It follows from the definition, the positive definity of and the lack of self-distance in terms of that the Wasserstein complexity is faithful: if and only if is the identity. The Wasserstein complexity is subbadditive under concatenation: . Indeed,
where the first inequality follows from the triangle inequality for the Wassertein divergence , the second inequality is due maximizing terms of a sum separately, and the last inequality follows from broadening the domain of the second maximum. From this it also follows directly that the Wasserstein complexity is subadditive under tensor products in the following sense:
We note that whether the Wassertein complexity (37) is convex or not is an open question.
5.2. Mean-field approximations of evolution equations
Among the many applications of classical optimal transport, one is particularly important for the development of quantum optimal transport theory. In [20], Dobrushin employed a special approximation based on a transport-related metric, the so-called Kantorovich-Rubinstein metric, to prove the uniqueness of the solution to the Vlasov equation. The Vlasov equations, which describe the limiting situation of weakly interacting particles with a large radius of interaction, are among the most frequently used kinetic equations in statistical mechanics. More recently, Golse, Paul, and Mouhot have extended Dobrushin’s approach to the quantum setting. Since Dobrushin used an optimal transport metric to compare -particle densities and their mean-field limits, it was a natural idea to define a transport-related quantity for the purpose of comparing quantum states [27]. The significance of the quantum counterpart of optimal transport (and the Wasserstein metric) in addressing various problems in quantum dynamics became even more evident later through their subsequent papers, see e.g. [9, 10, 27, 28, 31, 33].
5.3. The Luttinger model
The non-equilibrium dymamics of the Luttinger model, describing the low energy physics in Luttinger liquid is a field studied extensively. The question of how much the time evolved state described by differs from the initial state is often investigated. The paper [3] study the time evolution of Uhlmann fidelity (or the Loschmidt echo) which measures the overlap between the time evolved and the initial thermal equilibrium states, is evaluated for arbitrary initial temperatures and quench protocols. Instead of fidelity we can consider other dissimilarity measures between states to describe these effects. With zero initial temperature, i.e. the initial state is the pure ground state of the Hamiltonian, the Wasserstein distance is easiliy computable and can be used for further investigations.